Evaluate ∫0^{π/2α}αsinαxdx,α>0 by using Leibniz formula.`


1
Expert's answer
2021-08-25T09:28:02-0400

The Leibniz formula says

"\int_{a}^b f(x)dx=F(b)-F(a)"

We get


"\int_0^{π/2α}α\sinαxdx\\
=-\cos(\alpha*\pi/2\alpha)+\cos(0)=1"